How many cubic centimeters of ice cream can an ice cream cone (within the cone) hold if its height is 8 centimeters and the diameter of its base is 3 centimeters? Give solution accurate to the nearest hundredth. Include correct units with your solution.
Answer by checkley77(12844)

V=BH/3
V=3.14*3^2*8/3
V=3.14*9*8/3
V=3.14*72/2
V=3.14*36
V=113.04 CM^3.